Probability and Statistics I GSIA, Carnegie Mellon University

45-733, Spring 2002

Homework #1, due Thursday, January 24, 2002

Please do the following problems from the book:

pp 27-9  3,11,17,19
pp 89-91  1,3
pp 106-8  7,19
pp 117-127 39,47

Before you begin, please read section 3.7 of the book. I did not go over this in class because it is just an application of the material I did present. However, you are responsible for it and one of the problems does require it.

If you get stuck, it would be wise to try to work even-numbered problems similar to the ones assigned, since many of these have solutions in the back of the book.